A store owner has two different blends of coffee. Brand A sells for $10.50 a pound and Brand B sells for $5.75 a pound. The owner wants to create a 25-pound mixture of Brand A and B to sell for $8.22 a pound. How many pounds of each blend should he use?
Answer by stanbon(75887)

A store owner has two different blends of coffee. Brand A sells for $10.50 a pound and Brand B sells for $5.75 a pound. The owner wants to create a 25-pound mixture of Brand A and B to sell for $8.22 a pound. How many pounds of each blend should he use?
-----------------
Quantity Equation: A + B = 25 lb
Value Equation: 10.5A + 5.75B = 25*8.22
-----
Multiply thru the 1st Eq. by 1050
Multiply thru the 2nd Eq. by 100
------------
1050A + 1050B = 1050*25
1050A + 575B = 25*822
=========================
Subtract 2nd from 1st and solve for "B":
475B = 5700
B = 12 lb (amt. of Brand B needed in the mixture)
----
Since A+B = 25, A = 13 lb (amt. of Brand A needed in the mixture)
========================
Cheers,
Stan H.
===========